How many kg of a metal containing 20% nickel must be combined with 9 kg of pure nickel to form an alloy containing 92% nickel?
Answer by mananth(16946)

How many kg of a metal containing 20% nickel must be combined with 9 kg of pure nickel to form an alloy containing 92% nickel?
...
20% nickel ----------------- x kg
100% nickel-----------------9 kg
92% nickel...................(9+x) kg
...
20x+100*9= 92*(9+x)
20x+900=828+92x
92x-20x=900-828
72x=72
/72
x= 1kg of 20% nickel
m.ananth@hotmail.ca